The area of a rectangle is 250ft. You want the length to be 5ft longer than twice its width. What should the dimensions of the rectangle be.
Answer by tommyt3rd(5050)

A=lw=250
and
l=2w+5

so we can write
(2w+5)w=250
or
2w^2+5w-250=0
which has w=10 as one solution
therefore l=2(10)+5=25
so 10ft x 25ft


:)